# Measurement-induced phase transition in periodically driven free-fermionic systems

### v1

@@ III $2\times 2$ Toy model
While Fig. 3 focuses on the case with $\epsilon=0$, we now examine the dependence of $\overline{S_{E}(\gamma)}$ on $\epsilon$ in the high-frequency regime with $T=0.5$. Figure 4 shows that for $\epsilon=0$, $\overline{S_{E}(\gamma)}$ remains nearly zero for all $\gamma$, as also observed in Fig. 3. As $\epsilon$ increases, we observe a typical increase in $\overline{S_{E}(\gamma)}$. However, with further increases in $\gamma$, $\overline{S_{E}(\gamma)}$ approaches zero even for finite $\epsilon$, indicating a transition from depinning to pinning. From Fig. 4, it is also evident that the decay of $\overline{S_{E}(\gamma)}$ to zero occurs more rapidly for smaller $\epsilon$ compared to larger $\epsilon$. This suggests, indirectly, that the critical measurement strength $\gamma_{c}$ (if a measurement-induced transition exists) may decrease as $\epsilon$ decreases, favoring more the localized phase. In the following sections, we will now focus on the many-body Hamiltonian.
@@ IV Many-body system
In this section, we focus on the periodically driven many-body system described by the Hamiltonian in Eq. (1), which is continuously monitored. We initialize the system in a Néel state $|\psi_{0}\rangle=|1010\cdots\cdots 1010\rangle$ at time $t=0$ and let the system evolve according to Eq. (LABEL:nlse). In the absence of a drive, such a system undergoes a transition from the critical to the localized phase (depinning to pinning transition), which has been previously characterized using the scaling of entanglement entropy with system size Skinner et al. (2019); Minato et al. (2022b); Fava et al. (2023). In our work, we also use entanglement entropy as a diagnostic tool.
The entanglement between subsystem $A$ of length $l_{A}$ and the rest of the system is measured by the von Neumann entanglement entropy, which is defined as $S=-\text{tr}(\rho_{A}\ln\rho_{A})$ Amico et al. (2008), where the reduced density matrix $\rho_{A}$ is obtained by tracing over the degrees of freedom of the rest of the system (which we identify as $B$), i.e., $\rho_{A}=\text{tr}_{B}|\psi(t)\rangle\langle\psi(t)|$. Even in the continuously monitored driven system we are interested in, the evolution generators are quadratic, preserving the Gaussianity of the quantum trajectories. Hence, the entanglement entropy $S$ can be easily calculated using the eigenvalues $\lambda_{i}$ of the correlation matrix $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle$ as Peschel (2003); Peschel and Eisler (2009); Vidmar et al. (2017),
$S(l_{A},L)=-\sum_{i=1}^{l_{A}}\left[\lambda_{i}\ln\lambda_{i}+(1-\lambda_{i})\ln(1-\lambda_{i})\right].$
In our calculation, we take $l_{A}=L/2$, where $L$ is the total system size. Thus, the eigenvalues of the correlation matrix $D_{ij}$ are computed by restricting the indices up to $L/2$. After calculating the entanglement entropy for each trajectory, we average over different realizations, which we denote as $\overline{S}$. In the subsequent sections, we study the behavior of $\overline{S}$ in great detail for different frequency regimes.
@@ IV.1 High-frequency regime: Analytical prediction
It has been proposed recently that in the absence of the driving, the long-time dynamics of the continuously monitored fermionic Hamiltonian given by Eq. (1) could be theoretically described by an effective bosonized non-Hermitian sine-Gordon model  Buchhold et al. (2021),
@@
Here, $v=\partial_{k}E(k)|_{k=k_{F}},E(k)=2\cos k$ for free theory and $k_{F}$ is fermi momentum, $\hat{\phi}_{x}$ is the bosonic field operator with its conjugate field $\hat{\theta}_{x}$ obeying the commutation relation $[\partial_{x}{\hat{\theta}_{x}},\hat{\phi}_{x^{\prime}}]=-i\pi\delta(x-x^{\prime})$, and $\lambda=\gamma m^{2}$, where $m$ is $O(1)$, depends on normal ordering prescription. It has been shown recently that in a long-time steady state, the system comes to a dark state, which is defined as the zero energy eigenstate of the $H_{eff}$, i.e., $H_{eff}|\psi_{D}\rangle=0$ (or equivalently can be identified as an eigenstate of the evolution operator $U(T)=e^{-iH_{eff}T}$ with eigenvalues $1$, i.e., $U(T)|\psi_{D}\rangle=|\psi_{D}\rangle$). The steady-state entanglement entropy can be obtained from  Casini and Huerta (2009); Calabrese and Cardy (2004),
$S(l_{A})=\frac{1}{3}\langle\psi_{D}|\hat{\phi}_{x}\hat{\phi}_{x+l_{A}}|\psi_{D}\rangle.$ (8)
In the case of the square pulse driving protocol (as discussed in Sec. II), in a similar spirit, one can write down the effective Floquet evolution operator in terms of the non-Hermitian sine-Gordon Hamiltonian as
@@
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]},$
and in the high-frequency limit, the effective Floquet Hamiltonian $H_{eff}^{F}$ can be calculated using Magnus expansion as Bukov et al. (2015),
@@
where $\eta^{2}=1-\frac{4i\gamma}{\epsilon\pi v}$ Hence, one would expect that in such a high-frequency limit, the steady state of the driven continuously monitored system, as discussed in Sec. II, can be described by the dark state of the non-Hermitian sine-Gordon Hamiltonian Eq. (11) and steady state entanglement entropy can be obtained using Eq. (8), $|\psi_{D}\rangle$ correspond to the zero energy eigenstate of the Hamiltonian Eq. (11). Moreover, this also suggests that the entanglement dynamics in the low-frequency limit can effectively be mimicked by the continuously monitored no-drive dynamics of tight-binding Hamiltonian with effective hopping amplitude $J_{eff}=\epsilon/2$. It also automatically implies if in the no drive case, i.e., $\epsilon=2J$, the transition point is at $\gamma_{c}=\gamma_{0}$, for $\epsilon<2J$ (the drive is introduced), the transition point will be $\gamma_{c}=\epsilon\gamma_{0}/2J$, decays to zero linearly with decreasing $\epsilon$.
@@
$S=\frac{1}{3}C({\gamma})\ln(l_{A}),$ (12)
where $C({\gamma})$ is referred to as the effective central charge. We refer to this phase as a critical phase. One needs to find out renormalization flow equations to study the effect of the nonlinearity $\lambda$ and investigate the fate of the Gaussian fixed point. The Hamiltonian Eq. (11) can be brought in its action form using complex wick rotation $(x,t)\to({\tilde{\epsilon}}^{1/2}\tilde{\eta}^{1/2}x,i\tilde{\epsilon}^{-1/2}\tilde{\eta}^{-1/2}t)$, and the action $\mathcal{F}$ reads as,
@@
where $A$ is a positive number of order $O(1)$ determined by the propagator of the Gaussian theory. If increasing the RG iteration step $\lambda$ grows, it signifies that the non-linearity is relevant perturbation; on the other hand, if $\lambda$ decays, that will make the perturbation irrelevant, and the physics will still be governed by the Gaussian fixed point, implies the survival of the critical phase. In Fig. 5 we numerically solve the Eq. (14) with initial condition $|\lambda(s=0)|=0.1$ and $K(s=0,\epsilon)=\sqrt{\frac{2J}{\epsilon}{K^{2}(s=0,\epsilon=2J)}+1-\frac{2J}{\epsilon}}$, with $|K(s=0,\epsilon=2J)|=1.42\pi$. We choose the initial condition in such a way that for $\epsilon=2J$ (no-drive scenario), $\lambda$ decreases with increasing $s$, signifies RG flow towards the Gaussian fixed point, implying the critical phase. The question we ask is if we change $\epsilon$ (indirect way to introduce drive), how $\lambda$ will change with RG steps? We find that for $\epsilon=0.75J$ and $0.55J$, it decreases with increasing $s$. On the other hand, on decreasing $\epsilon$ further, the $\lambda$ starts growing with $s$, which makes the non-linear $\lambda$ term relevant. It signifies that physics is no longer governed by the Gaussian fixed point; instead, it is governed by a strong coupling fixed point, which is responsible for the area-law entanglement phase. This study also shows that decreasing $\epsilon$ promotes the area-law entanglement phase, and if one starts with an initial condition such that the no-drive scenario shows a critical phase, with decreasing $\epsilon$ one should be able to probe the critical-area law phase transition Alberton et al. (2021).
@@
It is straightforward to check that for a fixed non-zero $\gamma$, in the limit $\epsilon\to 0$, this correlation tends to zero, and the effective central charge $C(\gamma)$ (note that the leading order term in the entanglement entropy $S=\frac{C(\gamma)}{3}\ln l_{A}$) will also approach zero as $\sqrt{\frac{\epsilon}{\gamma}}$. This implies that in the case of $\epsilon=0$, for any finite $\gamma$, one would expect to observe the area-law phase. In a nutshell, we make three main predictions about the high-frequency limit drive: 1) the entanglement dynamics can be mimicked by the dynamics of no-drive continuously monitored free-fermionic chain with the hopping amplitude $J_{eff}=\epsilon/2$, 2) If one starts in critical-phase in the no-drive regime, i.e. $\epsilon=2J$, with decreasing $\epsilon$, one can be able to probe critical to area-law phase transition, 3) $\epsilon=0$ limit will not display entanglement transition, it will always show the area-law entanglement for any non-zero finite measurement strength, 4) the transition point $\gamma_{c}$ is expected to decay to zero linearly with decreasing $\epsilon$. Next, we will try to test our analytical prediction using numerical simulation on a finite-size tight-binding lattice.
@@
Figure 6: Stroboscopic growth of trajectory averaged entanglement entropy for $\gamma=0.05$ (upper panel) and $\gamma=0.20$ (lower panel). All plots are for $\epsilon=3J/4$ and $T=5$. The upper panel inset shows that in the high-frequency regime $T=0.5$, the numerical results are in agreement with no-drive results for $J_{eff}=3J/8$. Inset data is for $\gamma=0.05$ and $N=250$.
@@
Figure 7: Upper Panel: System size scaling of steady state trajectory averaged entanglement entropy $\overline{S(L)}$ for different $\gamma$. Lower Panel: Data collapse of entanglement entropy for different $L$ and $\gamma$ assuming the BKT type scaling ansatz with critical measurement strength $\gamma_{c}=0.086$. The results are for $T=5$ and $\epsilon=3J/4$. The inset in the lower panel shows the variation of the $\gamma_{c}$ (obtained from the data-collapse) for different $\epsilon$.
@@
Figure 8: Color variation corresponds to the different values of the effective central charge $C(\gamma)$ for $T=5$. Black Dots represent $\gamma_{c}$, and the dashed line is the best-fit line, identifying the phase transition.
@@ IV.2 Numerical results
In this section, we perform a numerical experiment on a finite-size lattice with periodic boundary conditions using the methodology described in Sec. II. First, we study the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial zero entanglement product state for different values of $\epsilon$ and measurement strengths $\gamma,$ for $T=5$. Figure. 6 shows the dynamics of $\overline{S(nT)}$ for $\epsilon=3J/4$. The upper panel shows for $\gamma=0.05$, the saturation value of the entanglement entropy increases with system size; on the other hand, the lower panel shows for $\gamma=0.2$, the steady state value of the average entanglement entropy hardly changes with system size. This is a signature that, with increasing measurement strength, an entanglement transition may occur, where for small values of $\gamma$, the steady-state entanglement increases with $L$, and for large $\gamma$, it does not scale with $L$. The inset of the upper panel shows the entanglement dynamics in high-frequency regime $T=0.5$ for $\gamma=0.05$ and $L=250$ (red-dotted line). As predicted in the previous section, we find that, indeed, for $T=0.5$, the entanglement dynamics can be mimicked by the continuously monitored no-drive scenario in a tight-binding lattice with an effective hopping strength $J_{eff}=\epsilon/2$, which is denoted by points in the inset figure. We further check the system size scaling of steady-state entanglement entropy for $\epsilon=3J/4$ in Fig. 7 (upper panel) for different $\gamma$ values. It shows clear evidence of a logarithmic-area law entanglement phase transition. To re-confirm that for small $\gamma$, the scaling is logarithmic (not volume law), we also perform an F-test (see Appendix. C). It has been shown recently that in the absence of driving, the measurement-induced critical to area-law phase transition with increasing $\gamma$ belongs to the BKT university class Szyniszewski et al. (2023); Alberton et al. (2021). The system size dependence of the entanglement entropy in the BKT universality class obeys the following scaling form in the vicinity of the critical $\gamma_{c}$ Harada and Kawashima (1997),

### v2

@@ III $2\times 2$ Toy model
While Fig. 3 focuses on the case with $\epsilon=0$, we now examine the dependence of $\overline{S_{E}(\gamma)}$ on $\epsilon$ in the high-frequency regime with $T=0.5$. Figure 4 shows that for $\epsilon=0$, $\overline{S_{E}(\gamma)}$ remains nearly zero for all $\gamma$, as also observed in Fig. 3. As $\epsilon$ increases, we observe a typical increase in $\overline{S_{E}(\gamma)}$. However, with further increases in $\gamma$, $\overline{S_{E}(\gamma)}$ approaches zero even for finite $\epsilon$, indicating a transition from depinning to pinning. From Fig. 4, it is also evident that the decay of $\overline{S_{E}(\gamma)}$ to zero occurs more rapidly for smaller $\epsilon$ compared to larger $\epsilon$. This suggests, indirectly, that the critical measurement strength $\gamma_{c}$ (if a measurement-induced transition exists) may decrease as $\epsilon$ decreases, favoring the localized phase more. In the following sections, we will now focus on the many-body Hamiltonian.
@@ IV Many-body system
In this section, we focus on the periodically driven many-body system described by the Hamiltonian in Eq. (1), which is continuously monitored. We initialize the system in a Néel state $|\psi_{0}\rangle=|1010\cdots\cdots 1010\rangle$ at time $t=0$ and let the system evolve according to Eq. (LABEL:nlse). A system undergoing a phase transition could be characterized using the scaling of entanglement entropy with system size Skinner et al. (2019a); Minato et al. (2022b); Fava et al. (2023). In our work, we also use entanglement entropy as a diagnostic tool.
The entanglement between subsystem $A$ of length $l$ and the rest of the system is measured by the von Neumann entanglement entropy, which is defined as $S=-\text{tr}(\rho_{A}\ln\rho_{A})$ Amico et al. (2008), where the reduced density matrix $\rho_{A}$ is obtained by tracing over the degrees of freedom of the rest of the system (which we identify as $B$), i.e., $\rho_{A}=\text{tr}_{B}|\psi(t)\rangle\langle\psi(t)|$. Even in the continuously monitored driven system we are interested in, the evolution generators are quadratic, preserving the Gaussianity of the quantum trajectories. Hence, the entanglement entropy $S$ can be easily calculated using the eigenvalues $\lambda_{i}$ of the correlation matrix $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle$ as Peschel (2003); Peschel and Eisler (2009); Vidmar et al. (2017),
$S(l,L)=-\sum_{i=1}^{l}\left[\lambda_{i}\ln\lambda_{i}+(1-\lambda_{i})\ln(1-\lambda_{i})\right].$ (7)
In volume law phase, entanglement entropy increases linearly with subsystem size $l$; on the other hand, in area law phase, it is independent of the subsystem size. CFT with periodic boundary predicts the growth of the entanglement entropy in the critical phase as  Alberton et al. (2021),
$S(l,L)=\frac{c(\gamma)}{3}\ln(\frac{L}{\pi}\sin\frac{\pi l}{L})+S_{0}(\gamma),$ (8)
where $c(\gamma)$ is the effective central charge and $S_{0}(\gamma)$ is the residual entropy. After calculating the entanglement entropy for each trajectory, we average over different realizations, which we denote as $\overline{S}$. In the subsequent sections, we study the behavior of $\overline{S}$ in great detail for different frequency regimes.
@@ IV.1 High-frequency regime: Analytical prediction
It has been proposed that in the absence of the driving, the long-time dynamics of the continuously monitored fermionic Hamiltonian given by Eq. (1) could be theoretically described by an effective bosonized non-Hermitian sine-Gordon model  Buchhold et al. (2021),
@@
Here, $v=\partial_{k}E(k)|_{k=k_{F}},E(k)=2\cos k$ for free theory and $k_{F}$ is fermi momentum, $\hat{\phi}_{x}$ is the bosonic field operator with its conjugate field $\hat{\theta}_{x}$ obeying the commutation relation $[\partial_{x}{\hat{\theta}_{x}},\hat{\phi}_{x^{\prime}}]=-i\pi\delta(x-x^{\prime})$, and $\lambda=\gamma m^{2}$, where $m$ is $O(1)$, depends on normal ordering prescription. It has been shown that in a long-time steady state, the system comes to a dark state, which is defined as the zero energy eigenstate of the $H_{eff}$, i.e., $H_{eff}|\psi_{D}\rangle=0$ (or equivalently can be identified as an eigenstate of the evolution operator $U(T)=e^{-iH_{eff}T}$ with eigenvalues $1$, i.e., $U(T)|\psi_{D}\rangle=|\psi_{D}\rangle$). The steady-state entanglement entropy can be obtained from  Casini and Huerta (2009); Calabrese and Cardy (2004),
$S(l)=\frac{1}{3}\langle\psi_{D}|\hat{\phi}_{x}\hat{\phi}_{x+l}|\psi_{D}\rangle.$ (10)
Note that this analysis has been questioned recently with the growing belief of the absence of measurement-induced phase transition in these models, and the analysis predicts a critical to area law phase transition in the no-drive scenario. However, since this study still captures the essential features of the measurement-induced transition and can be extended to the driven system—at least in the extreme high-frequency limit—it motivates us to carry out this analysis to develop some intuition about the driven case.
In case of the square pulse driving protocol (as discussed in Sec. II), in a similar spirit, one can write down the effective Floquet evolution operator in terms of the non-Hermitian sine-Gordon Hamiltonian as
@@
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]}.$
Construction of Floquet operator where a non-hermitian Hamiltonian governs the dynamics has been studied in Ref. Zhou (2024); Chakrabarty and Datta (2024b). In the high-frequency limit, the effective Floquet Hamiltonian $H_{eff}^{F}$ can be calculated using Magnus expansion as Bukov et al. (2015),
@@
where $\eta^{2}=1-\frac{4i\gamma}{\epsilon\pi v}$ Hence, one would expect that in such a high-frequency limit, the steady state of the driven continuously monitored system, as discussed in Sec. II, can be described by the dark state of the non-Hermitian sine-Gordon Hamiltonian Eq. (13) and steady state entanglement entropy can be obtained using Eq. (10), $|\psi_{D}\rangle$ correspond to the zero energy eigenstate of the Hamiltonian Eq. (13). Moreover, this also suggests that the entanglement dynamics in the high-frequency limit can effectively be mimicked by the continuously monitored no-drive dynamics of tight-binding Hamiltonian with effective hopping amplitude $J_{eff}=\epsilon/2$. We checked numerically in Section. IV.2 in Fig. 6 upper panel inset, taking a very small time period $T=0.5$ for a given $\epsilon$ and the dynamics is completely matching with $J=\epsilon/2$ model. It also automatically implies if in the no drive case, i.e., $\epsilon=2J$, the transition point is at $\gamma_{c}=\gamma_{0}$ (which is predicted by our RG analysis), for $\epsilon<2J$ (the drive is introduced), the transition point will be $\gamma_{c}\simeq\epsilon\gamma_{0}/2J$, decays to zero linearly with decreasing $\epsilon$.
@@
$S=\frac{1}{3}c({\gamma})\ln(l),$ (14)
where $c({\gamma})$ is referred to as the effective central charge. We refer to this phase as a critical phase. One needs to find out the renormalization flow equations to study the effect of the nonlinearity $\lambda$ and investigate the fate of the Gaussian fixed point. The Hamiltonian Eq. (13) can be brought in its action form using complex wick rotation $(x,t)\to({\tilde{\epsilon}}^{1/2}{\eta}^{1/2}x,i\tilde{\epsilon}^{-1/2}{\eta}^{-1/2}t)$, and the action $\mathcal{F}$ reads as,
@@
where $A$ is a positive number of order $O(1)$ determined by the propagator of the Gaussian theory. If increasing the RG iteration step $\lambda$ grows, it signifies that the non-linearity is a relevant perturbation; on the other hand, if $\lambda$ decays, that will make the perturbation irrelevant, and the physics will still be governed by the Gaussian fixed point, implying the survival of the critical phase. In Fig. 5 we numerically solve the Eq. (16) with initial condition $|\lambda(s=0)|=0.1$ and $K(s=0,\epsilon)=\sqrt{\frac{2J}{\epsilon}{K^{2}(s=0,\epsilon=2J)}+1-\frac{2J}{\epsilon}}$, with $|K(s=0,\epsilon=2J)|=1.42\pi$. These initial conditions are not special apart from the fact that for such initial conditions in the $\epsilon=2J$ limit (no-drive scenario), $\lambda$ always decreases with increasing $s$, signifying RG flow towards the Gaussian fixed point, implying the critical phase. The question we ask is: if we change $\epsilon$ (an indirect way to introduce drive), how will $\lambda$ change with RG steps? We see, on decreasing $\epsilon$, the $\lambda$ starts growing with $s$, which makes the non-linear $\lambda$ term relevant, while for $\epsilon=0.5J$ we get that the flow is marginal for our chosen initial condition. Also, we have checked that with increasing the initial value of the parameter $|K(\epsilon=2J)|$, the marginal $\epsilon$ value also increases. It signifies that physics is no longer governed by the Gaussian fixed point; instead, it is governed by a strong coupling fixed point, which is responsible for the area-law entanglement phase. This study also shows that decreasing $\epsilon$ promotes the area-law entanglement phase, and if one starts with an initial condition such that the no-drive scenario shows a critical phase, with decreasing $\epsilon$, one should be able to probe the critical-area law phase transition Alberton et al. (2021).
@@
It is straightforward to check that for a fixed non-zero $\gamma$, in the limit $\epsilon\to 0$, this correlation tends to zero, and the effective central charge $c(\gamma)$ (note that the leading order term in the entanglement entropy $S=\frac{c(\gamma)}{3}\ln l$) will also approach zero as $\sqrt{\frac{\epsilon}{\gamma}}$. This implies that in the case of $\epsilon=0$, for any finite $\gamma$, one would expect to observe the area-law phase. In a nutshell, we make four main predictions based on this RG analysis about the high-frequency limit drive: 1) the entanglement dynamics can be mimicked by the dynamics of no-drive continuously monitored free-fermionic chain with the hopping amplitude $J_{eff}=\epsilon/2$, 2) If one starts in critical-phase in the no-drive regime, i.e. $\epsilon=2J$, with decreasing $\epsilon$, one can be able to probe critical to area-law phase transition, 3) $\epsilon=0$ limit will not display entanglement transition, it will always show the area-law entanglement for any non-zero finite measurement strength, 4) the transition point $\gamma_{c}$ is expected to decay to zero linearly with decreasing $\epsilon$ in the high-frequency region.
In this context, we would like to mention that very recently, a modified version of the RG analysis on sine-Gordon type models has been proposed after refining their previous analysis, in Ref. Müller et al. (2025), which shows the critical to area law transition point shifts towards zero in the case of a non-interacting monitored system in the absence of drive. However, the symmetries and measurement protocols of these continuum models are not exactly identical to our lattice setup. In monitored systems, energy conservation is broken, and measurements can populate fermions across the entire band. Hence, there is no guarantee that any such models (that include Hamiltonian Eq. (9) too) can describe the dynamics of monitored fermions on a tight-binding lattice. Also, the prediction of Ref. Müller et al. (2025) of a logarithmic to area law phase transition in the interacting case, contradicts the other recent works Poboiko et al. (2025); Xing et al. (2024); Guo et al. (2024), where they observed a volume-law to area-law phase transition with increasing measurement rate. Overall, it implies that it is a topic of debate what is the appropriate effective continuum model and the corresponding correct RG equations that can describe the high-frequency limit of a driven continuously monitored system, but in our study, it helps us to build an intuition which we are going to test in the subsequent section with the finite-sized numerics.
@@
Figure 6: Stroboscopic growth of trajectory averaged entanglement entropy for $\gamma=0.05$ (upper panel) and $\gamma=0.20$ (lower panel). All plots are for $\epsilon=3J/4$ and $T=5$. The upper panel inset shows that in the high-frequency regime $T=0.5$, the numerical results are in agreement with no-drive results for $J_{eff}=3J/8$. Inset data is for $\gamma=0.05$ and $L=250$.
@@
Figure 7: Upper Panel: System size scaling of steady state trajectory averaged entanglement entropy $\overline{S(L)}$ for different $\gamma$. Lower Panel: Data collapse of entanglement entropy for different $L$ and $\gamma$ assuming the BKT type scaling ansatz with critical measurement strength $\gamma_{c}=0.08$. The results are for $T=5$ and $\epsilon=3J/4$. The inset in the lower panel shows the variation of the $\gamma_{c}$ (obtained from the data-collapse) for different $\epsilon$.
@@
Figure 8: Color variation corresponds to the different values of the effective central charge $c(\gamma)$ for $T=5$. Black Dots represent $\gamma_{c}$ for finite-size system.
@@ IV.2 Numerical results
In this section, we perform a numerical experiment on a finite-size lattice with periodic boundary conditions using the methodology described in Sec. II. We follow the following strategies; first, we will pretend the BKT transition takes place for the driven system, and compute critical measurement strength $\gamma_{c}$ using appropriate BKT ansatz, then we compute the local effective central charge (19) as proposed in Ref. Starchl et al. (2024a), and investigate whether for $\gamma<\gamma_{c}$ (obtained from scaling analysis) the local effective central charge start decreasing with increasing system size. If we don’t find such a signature, then we would conclude that the critical phase is stable.
@@
As a beginning, we study the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial zero entanglement product state for different values of $\epsilon$ and measurement strengths $\gamma,$ for $T=5$. Figure. 6 shows the dynamics of $\overline{S(nT)}$ for $\epsilon=3J/4$. The upper panel shows that for $\gamma=0.05$, the saturation value of the entanglement entropy increases with system size; on the other hand, the lower panel shows that for $\gamma=0.2$, the steady state value of the average entanglement entropy hardly changes with system size. This is a signature that, with increasing measurement strength, an entanglement transition may occur, where for small values of $\gamma$, the steady-state entanglement increases with $L$, and for large $\gamma$, it does not scale with $L$. The inset of the upper panel shows the entanglement dynamics in the high-frequency regime $T=0.5$ for $\gamma=0.05$ and $L=250$ (red-dotted line). As predicted in the previous section, we find that, indeed, for $T=0.5$, the entanglement dynamics can be mimicked by the continuously monitored no-drive scenario in a tight-binding lattice with an effective hopping strength $J_{eff}=\epsilon/2$, which is denoted by points in the inset figure. Given, the results for $T<1$ are almost indistinguishable from the no-drive results, at least for our accessible system size, hence, we consider $T=5>1$ for most of our numerics. The no-drive results are separately presented in the Appendix. D.
Next, we want to see how entanglement entropy scales with the subsystem sizes, and want to predict the critical $\gamma_{c}$(if it exists) at which an entanglement phase transition happens. For this, we opt for two methods. In one case (methodology 1), we vary the total system size $L$ and always take the subsystem size as $l=L/2$. In the other method (Methodology 2), we take a large system size $L$, and calculate the entanglement entropy for different subsystem sizes $l$.
Methodology 1:
We check the system size scaling of steady-state entanglement entropy for $\epsilon=3J/4$ in Fig. 7 (upper panel) for different $\gamma$ values. It shows evidence of a logarithmic-area law entanglement phase transition. To reconfirm that for small $\gamma$, the scaling is logarithmic (not volume law), we also perform an F-test (see Appendix. C). A plethora of studies in the last few years claim that, in the absence of driving, the measurement-induced critical-to-area-law phase transition with increasing $\gamma$ belongs to the BKT universality class Szyniszewski et al. (2023); Alberton et al. (2021); Müller et al. (2022); Minato et al. (2022a). However, the existence of such a transition in the thermodynamic limit has been questioned very recently; it has been proposed that even for small $\gamma$, the steady-state entanglement entropy will show area-law scaling Poboiko et al. (2023). At least for numerically accessible system sizes and for $\epsilon=3J/4$, we do not observe such signatures; the data mostly support the logarithmic scaling of the steady-state entanglement entropy with system size. Hence, we fit our data with the BKT scaling ansatz. The system size dependence of the entanglement entropy in the BKT universality class obeys the following scaling form in the vicinity of the critical $\gamma_{c}$ Harada and Kawashima (1997),